X is a discrete random variable whose possible values are all integers. It is known that Fx(3) = .5, Fx(4) = .65, Fx(5) = .75 and Fx(6) = .95. Find Pr[X > 4].

A. .25

B. .35

C. .5

D. .75

E. .85
Answer #1

P(X > 4) = 1 - P(X < 4) = 1 - F(4) = 1 - 0.65 = 0.35

Option-B) 0.35